# CUBIC DIOPHANTINE EQUATION SATISFIABILITY IS $\Sigma^{0}_{1}$-COMPLETE

### v6

@@ 4.9 Compilation and Degree Control
Definition 4.29 (Cubic system compiler).
@@
$\mathtt{emit\_cubic\_system}:\mathsf{Proof}\times\mathsf{Form}\to\mathsf{List}(\mathsf{Polynomial}),$ (4.107)
@@
Definition 4.30 (Single-equation compiler).
@@
$\mathtt{emit\_single\_cubic}:\mathsf{Proof}\times\mathsf{Form}\to(\mathsf{Polynomial}\times\mathsf{Polynomial})$ (4.108)
@@
Theorem 4.31 (Certified cubic reduction).
@@
$\mathtt{emit\_cubic\_system}(p,\varphi)$ (4.109)
@@
$\mathrm{check}(p,\varphi)=\mathrm{true}\Leftrightarrow\exists a.\;\mathtt{all\_zero}\bigl(\mathtt{emit\_cubic\_system}(p,\varphi),\;a\bigr).$ (4.110)
@@
$\mathtt{check\_iff\_emit\_cubic\_all\_zero}.$ (4.111)
By mechanization. In Rocq: degree bounds by structural induction on emitted constraints ($\mathtt{Constraints\_Assembly}$), and correctness by the checker–constraint equivalence ($\mathtt{Constraints\_MP}$). The reduction nucleus is $\mathtt{compiler\_primitive}$ in $\mathtt{R16}$. ∎
Remark.
The classical monomial-introduction construction of Jones [1982] remains visible as an alternative route from degree-$6$ back to degree-$3$ systems, but is not used in the present argument.
@@ 5.1 On Local Reflection
Intuition.
Every nonexistence result in this section proceeds by modus tollens: assume the object exists and derive $\bot$. This is not a deficiency in our methodology but reflects the intrinsic difficulty of proving a negative statement. Consider, for example, how under the Brouwer–Heyting–Kolmogorov interpretation [Troelstra and van Dalen, 1988], absurdity admits no canonical realization. Accordingly, there is no direct witness of impossibility. To establish that a decider or any promissory procedure does not exist, the only known route is to show that any hypothetical witness yields a contradiction under suitable assumptions. Proving impossibility therefore amounts to demonstrating that the assumption is untenable within a consistency-preserving environment.
In our case, structural constraints determine the specific form of the transformation. Once the certified compiler of Theorem 4.31 reliably produces a system of cubic constraints for each proof target, while a two-channel aggregation (Lemma 4.25) witnesses the map from theoremhood to a single cubic equation without change in degree, we argue via an induced machine that could observe itself: the compiler $f$, composed with the aggregation, is itself primitive recursive and hence computable, such that any hypothetical total correct decider $D$ for $\mathsf{CubicSat}$ would compose with $f$ to decide $\mathsf{Thm}$. It is the existence of this induced oracle machine, not a specific impossible equation, that is refuted.
Ansatz 5.1 (Argument).
@@
    Internal collapse. If correctness is provable inside $\mathsf{RA}$ on the compiler image, one derives $\Sigma_{1}$-reflection and contradicts Gödel’s 2nd.
@@
Intuition.
A “total machine” $\mathcal{M}$ that correctly identifies all unsolvable instances—reporting $M(e)=0$ whenever $\neg\mathsf{DiophEqSat}(e)$—is, together with $\Sigma_{1}$-enumeration of solutions, a total decider for $\mathsf{DiophEqSat}$. Detecting unsolvability and deciding solvability mutually imply each other, such that internally certifying inside $\mathsf{RA}$ that $\mathcal{M}$ correctly identifies unsolvable instances enables a $\Pi_{1}$-soundness schema on the cubic predicate:
$\operatorname{Prov}_{\mathsf{RA}}\bigl(\ulcorner\neg\mathsf{DiophEqSat}(\bar{e})\urcorner\bigr)\;\to\;\neg\mathsf{DiophEqSat}(e).$ (5.112)
This is a form of Local Reflection Principle [Feferman, 1960]. We observe how $\mathcal{M}$ escalates to $\mathrm{Con}(\mathsf{RA})$, which is false in $\mathsf{RA}$ by Gödel’s 2nd. Any internally certified total primitive procedure for detecting unsolvability forces $\mathsf{RA}$ to prove its own consistency: a contradiction.
@@ 5.2 Fixed Point Logic
Theorem 5.2 (No total correct single-cubic decider).
@@
$\forall u\;\bigl(\mathsf{Thm}(u)\leftrightarrow\mathsf{CubicSat}(f(u))\bigr),$ (5.113)
where $\mathsf{CubicSat}$ is satisfiability of a single degree-$\leq 3$ Diophantine equation over $\mathbb{N}$ (the system emitted by the compiler is collapsed to one equation by Fibonacci-banded aggregation, Lemma 4.25).
Since $\mathsf{CubicSat}$ is already the satisfiability of a single cubic equation, write $\mathsf{CubicEqSat}:=\mathsf{CubicSat}$. Then there is no total Turing Machine $D$ such that
$\forall s\;\bigl(D(s)=1\leftrightarrow\mathsf{CubicEqSat}(s)\bigr).$ (5.114)
@@
By Fixed point logic in steps. We consult Boolos et al. [2012]; Kleene [1952]
@@
    Fix the primitive recursive map $f$ witnessing (5.113). Since $\mathsf{CubicSat}$ is itself satisfiability of a single cubic equation (the compiler-emitted system is collapsed by Fibonacci-banded aggregation, Lemma 4.25), $f$ already maps sentence codes to single cubic equation codes. Define h(u):=f(u),h(u)\;:=\;f(u), observe that $h$ is primitive recursive and satisfies
    $\forall u\;\bigl(\mathsf{Thm}(u)\leftrightarrow\mathsf{CubicEqSat}(h(u))\bigr).$ (5.115)
@@
    We assume a total correct single-cubic decider. For contradiction, there exists a total Turing Machine $D$ satisfying (5.114). Let $\mathsf{Out}_{D}(s)$ be the arithmetization:
@@
    $\varphi(x)\;\equiv\;\neg\,\mathsf{Out}_{D}(h(x)).$ (5.117)
    By the Diagonal Lemma (Theorem 3.14), there exists a sentence $u_{D}$ such that
    $\mathsf{RA}\;\vdash\;u_{D}\leftrightarrow\neg\,\mathsf{Out}_{D}\bigl(h(\ulcorner u_{D}\urcorner)\bigr).$ (5.118)
@@
    Case 1. $\mathsf{Out}_{D}(s_{D})$, then by (5.118) $\neg\,\mathsf{Thm}(\ulcorner u_{D}\urcorner)$, but by (5.116) and correctness (5.114), $\mathsf{CubicEqSat}(s_{D})$ holds, (5.115) yields $\mathsf{Thm}(\ulcorner u_{D}\urcorner)$, contradiction.
    Case 2. $\neg\,\mathsf{Out}_{D}(s_{D})$, then by (5.118), $\mathsf{Thm}(\ulcorner u_{D}\urcorner)$, hence (5.115) gives $\mathsf{CubicEqSat}(s_{D})$, and by correctness (5.114) plus totality (5.116) we obtain $\mathsf{Out}_{D}(s_{D})$, contradiction.
Both cases contradict, so the assumption (5.114) is false. Classically, no correct Turing Machine decides $\mathsf{CubicEqSat}$. ∎
Intuition.
The only $\mathsf{RA}$ argument required for the above is the fixed-point statement (5.118), obtained from the Diagonal Lemma for $\mathsf{RA}$ and the representability of the primitive recursive map $h$. The contradiction itself is external: it uses the totality/correctness assumption about $D$ to identify $\mathsf{Out}_{D}(s)$ with $D(s)=1$, and uses the reduction
$\mathsf{Thm}(u)\leftrightarrow\mathsf{CubicEqSat}(h(u))$ (5.120)
@@
Corollary 5.3 (Undecidability for single Diophantine equations).
@@
$\mathsf{DiophEqSat}(e)\;\equiv\;\exists\,\vec{x}\in\mathbb{N}^{k}\;P_{e}(\vec{x})=0,$ (5.121)
where $e$ codes a single polynomial $P_{e}\in\mathbb{Z}[x_{1},\dots,x_{k}]$ (with $k$ part of the code), is $\Sigma^{0}_{1}$-complete and undecidable. In particular, the set of satisfiable single cubic equations over $\mathbb{N}$ (with arity encoded in the instance) is undecidable; over $\mathbb{Z}$ the corresponding degree bound is $6$ by difference encoding.
Since $\mathsf{DiophEqSat}\in\Sigma^{0}_{1}$, hardness and membership together yield $\Sigma^{0}_{1}$-completeness, and undecidability follows. ∎
Remark.
@@ 5.3 Certification and Reflection
Definition 5.4 (Internal predicates).
We use the following arithmetized predicates (all are primitive recursive and representable in $\mathsf{RA}$):
@@
  • $\operatorname{Dioph}_{3}(e)$: the cubic polynomial coded by $e$ has a zero over $\mathbb{N}$ (same as $\mathsf{CubicEqSat}$).
  • $\operatorname{HaltOut}_{\mathcal{M}}(s,b)$: machine $M$ halts on input $s$ with output bit $b$.
@@
Intuition.
$\operatorname{Prov}_{\mathsf{RA}}(u)$$u$ (code)Compiler fff(u)f(u)Decider $\mathcal{M}$$\operatorname{tr}$ (bridge) Internal reasoning in $\mathsf{RA}$ $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}\ \Longrightarrow\ \mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$ $\mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})\ \Longrightarrow\ \mathsf{RA}\vdash\mathrm{Con}(\mathsf{RA})\Longrightarrow\bot$
Figure 2: Internal certification loop. If $\mathsf{RA}$ could witness a machine $\mathcal{M}$ that decides the compiler image ($\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$), then (via the bridge $\operatorname{tr}$ and representability of $f$) provability transports to $\Sigma_{1}$-truth on that image, yielding $\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$, hence $\mathrm{Con}(\mathsf{RA})$, contradicting Gödel’s 2nd.
As said, certification of a cubic decider inside $\mathsf{RA}$ would not merely refute the decider—it would equip $\mathsf{RA}$ with the means to prove its own consistency by a Reflection Principle. The provability bridge transports $\mathsf{RA}$-proofs through the compiler $f$; the certified decision specification $\mathrm{Dec}_{\mathcal{M}}$ then closes the reflection loop (Figure 2).
Definition 5.5 (Internal decision specification).
Let $\operatorname{Sat}_{\Sigma_{1}}(u)$ be the standard $\Sigma_{1}$-satisfaction predicate and $f$ the primitive recursive compiler of Theorem 1.8. For a Turing Machine $\mathcal{M}$, define $\mathrm{Dec}_{\mathcal{M}}$ as
$\forall u\;\Big(\operatorname{Sigma1Sent}(u)\rightarrow\big(\operatorname{HaltOut}_{\mathcal{M}}(f(u),1)\leftrightarrow\operatorname{Dioph}_{3}(f(u))\big)\Big).$ (5.122)
@@
Theorem 5.6 (Provable decision on the compiler image yields $\Sigma_{1}$-reflection).
@@
$\mathsf{RA}\vdash\;\forall u\;\Big(\operatorname{Sigma1Sent}(u)\rightarrow\big(\operatorname{Sat}_{\Sigma_{1}}(u)\leftrightarrow\operatorname{Dioph}_{3}(f(u))\big)\Big).$ (5.123)
@@
Since $f$ is primitive recursive, its graph is $\mathsf{RA}$-representable (Proposition 1.10). Internalizing the compilation step yields a provability bridge: for each $\Sigma_{1}$-sentence code $u$,
@@
$\mathsf{RA}\vdash\;\operatorname{Prov}_{\mathsf{RA}}(u)\rightarrow\operatorname{Prov}_{\mathsf{RA}}\bigl(\ulcorner\operatorname{Dioph}_{3}(f(u))\urcorner\bigr).$ (5.124)
@@
Combining with $\mathrm{Dec}_{\mathcal{M}}$ and (5.123) yields $\operatorname{Prov}_{\mathsf{RA}}(u)\rightarrow\operatorname{Sat}_{\Sigma_{1}}(u)$, i.e. $\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$. ∎
Corollary 5.7 (Incompleteness Discharge).
@@
Assume $\mathsf{RA}$ is consistent with $\operatorname{Prov}_{\mathsf{RA}}$ satisfying the Hilbert–Bernays–Löb conditions (D1)–(D3) of Theorem 3.13. Then no Turing Machine $\mathcal{M}$ satisfies $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$: Theorem 5.6 would give $\mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$, hence $\mathsf{RA}\vdash\mathrm{Con}(\mathsf{RA})$—contradicting Gödel’s 2nd Gödel [1931]; Feferman [1960]; Boolos et al. [2012]. If $\mathrm{Con}(\mathsf{RA})$ then $\nexists\mathcal{M}$, if $\lnot\mathrm{Con}(\mathsf{RA})$ then also $\nexists\mathcal{M}$.
The argument has three strictly separated tiers:
  1. Tier 1.

### v7

@@ 4.9 Compilation and Degree Control
Definition 4.9.2 (Cubic system compiler).
@@
$\mathtt{emit\_cubic\_system}:\mathsf{Proof}\times\mathsf{Form}\to\mathsf{List}(\mathsf{Polynomial}),$ (130)
@@
Definition 4.9.3 (Single-equation compiler).
@@
$\mathtt{emit\_single\_cubic}:\mathsf{Proof}\times\mathsf{Form}\to(\mathsf{Polynomial}\times\mathsf{Polynomial})$ (131)
@@
Mechanization 4.9.4 (Certified cubic reduction).
@@
$\mathtt{emit\_cubic\_system}(p,\varphi)$ (132)
@@
$\mathrm{check}(p,\varphi)=\mathrm{true}\Leftrightarrow\exists a.\;\mathtt{all\_zero}\bigl(\mathtt{emit\_cubic\_system}(p,\varphi),\;a\bigr).$ (133)
@@
$\mathtt{check\_iff\_emit\_cubic\_all\_zero}.$ (134)
@@ 5.1 On Local Reflection
Remark 5.1.1.
The compiler $f$ of Theorem 4.9.4, composed with two-channel aggregation (Lemma 4.8.5), is primitive recursive. Any hypothetical total correct decider $D$ for $\mathsf{CubicSat}$ would therefore compose with $f$ to decide $\mathsf{Thm}$. It is the existence of this induced machine, not a specific impossible equation, that is refuted.
Ansatz 5.1.2 (Argument).
@@
    Internal collapse. If correctness is provable inside $\mathsf{RA}$ on the compiler image, one derives $\Sigma_{1}$-reflection and contradicts Gödel’s 2nd by modus tollens.
@@
Remark 5.1.3.
A total machine $\mathcal{M}$ correctly identifying all unsolvable instances would, together with $\Sigma_{1}$-enumeration, yield a total decider for $\mathsf{DiophEqSat}$. Internally certifying $\mathcal{M}$ inside $\mathsf{RA}$ enables a $\Pi_{1}$-soundness schema:
$\operatorname{Prov}_{\mathsf{RA}}\bigl(\ulcorner\neg\mathsf{DiophEqSat}(\bar{e})\urcorner\bigr)\;\to\;\neg\mathsf{DiophEqSat}(e),$ (135)
a local reflection principle [Feferman, 1960] admitting $\mathsf{RA}\vdash\mathrm{Con}(\mathsf{RA})$, hence absurdum by contradicting Gödel’s 2nd.
@@ 5.2 Fixed Point Logic
Theorem 5.2.1 (No total correct single-cubic decider).
@@
$\forall u\;\bigl(\mathsf{Thm}(u)\leftrightarrow\mathsf{CubicSat}(f(u))\bigr),$ (136)
where $\mathsf{CubicSat}$ is satisfiability of a single degree-$\leq 3$ Diophantine equation over $\mathbb{N}$ (the system emitted by the compiler is collapsed to one equation by Fibonacci-banded aggregation, Lemma 4.8.5).
Since $\mathsf{CubicSat}$ is already the satisfiability of a single cubic equation, it coincides with the canonical predicate $\mathsf{CubicEqSat}$ fixed in Equation (28). Then there is no total Turing Machine $D$ such that
$\forall s\;\bigl(D(s)=1\leftrightarrow\mathsf{CubicEqSat}(s)\bigr).$ (137)
@@
By diagonal in steps. We consult Boolos et al. [2012]; Kleene [1952].
@@
    Fix the primitive recursive map $f$ witnessing (136). Since $\mathsf{CubicSat}$ is itself satisfiability of a single cubic equation (the compiler-emitted system is collapsed by Fibonacci-banded aggregation, Lemma 4.8.5), $f$ already maps sentence codes to single cubic equation codes. Define h(u):=f(u),h(u)\;:=\;f(u), observe that $h$ is primitive recursive and satisfies
    $\forall u\;\bigl(\mathsf{Thm}(u)\leftrightarrow\mathsf{CubicEqSat}(h(u))\bigr).$ (138)
@@
    We assume a total correct single-cubic decider. For contradiction, there exists a total Turing Machine $D$ satisfying (137). Let $\mathsf{Out}_{D}(s)$ be the arithmetization:
@@
    $\varphi(x)\;\equiv\;\neg\,\mathsf{Out}_{D}(h(x)).$ (140)
    By the Diagonal Lemma (Theorem 3.5.2), there exists a sentence $u_{D}$ such that
    $\mathsf{RA}\;\vdash\;u_{D}\leftrightarrow\neg\,\mathsf{Out}_{D}\bigl(h(\ulcorner u_{D}\urcorner)\bigr).$ (141)
@@
    Case 1. $\mathsf{Out}_{D}(s_{D})$, then by (141) $\neg\,\mathsf{Thm}(\ulcorner u_{D}\urcorner)$, but by (139) and correctness (137), $\mathsf{CubicEqSat}(s_{D})$ holds, (138) yields $\mathsf{Thm}(\ulcorner u_{D}\urcorner)$, contradiction.
    Case 2. $\neg\,\mathsf{Out}_{D}(s_{D})$, then by (141), $\mathsf{Thm}(\ulcorner u_{D}\urcorner)$, hence (138) gives $\mathsf{CubicEqSat}(s_{D})$, and by correctness (137) plus totality (139) we obtain $\mathsf{Out}_{D}(s_{D})$, contradiction.
Both cases contradict, so the assumption (137) is false. Classically, no correct Turing Machine decides $\mathsf{CubicEqSat}$. ∎
Remark 5.2.2.
The only $\mathsf{RA}$ argument required for the above is the fixed-point statement (141), obtained from the Diagonal Lemma for $\mathsf{RA}$ and the representability of the primitive recursive map $h$. The contradiction itself is external: it uses the totality/correctness assumption about $D$ to identify $\mathsf{Out}_{D}(s)$ with $D(s)=1$, and uses the reduction
$\mathsf{Thm}(u)\leftrightarrow\mathsf{CubicEqSat}(h(u))$ (143)
@@
Corollary 5.2.3 (Undecidability for single Diophantine equations).
@@
$\mathsf{DiophEqSat}(e)\;\equiv\;\exists\,\vec{x}\in\mathbb{N}^{k}\;P_{e}(\vec{x})=0,$ (144)
where $e$ codes a single polynomial $P_{e}\in\mathbb{Z}[x_{1},\cdots,x_{k}]$ (with $k$ part of the code), is $\Sigma^{0}_{1}$-complete and undecidable. In particular, the set of satisfiable single cubic equations over $\mathbb{N}$ (with arity encoded in the instance) is undecidable; over $\mathbb{Z}$ the corresponding degree bound is $6$ by difference encoding.
Immediate. Since $\mathsf{DiophEqSat}\in\Sigma^{0}_{1}$, hardness and membership together yield $\Sigma^{0}_{1}$-completeness, and undecidability follows. ∎
Remark 5.2.4.
@@ 5.3 Certification and Reflection
Definition 5.3.1 (Internal predicates).
[…]
Figure 2: Internal certification loop (Magenta). If $\mathsf{RA}$ could witness a machine $\mathcal{M}$ that decides the compiler image ($\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$), via the bridge $\operatorname{tr}$ (Green) and representability of $f$, provability transports to $\Sigma_{1}$-truth on that image, yielding $\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$, hence impossibility under consistency.
Consider the following arithmetized predicates (all are primitive recursive and representable in $\mathsf{RA}$):
  1. (i)
@@
  3. (iii)
    $\operatorname{Dioph}_{3}(e)$: the cubic polynomial coded by $e$ has a zero over $\mathbb{N}$ (the canonical predicate $\mathsf{CubicEqSat}$ of Equation 28; the pair surface used by the kernel is equivalent by Lemma 2.2.3).
  4. (iv)
    $\operatorname{HaltOut}_{\mathcal{M}}(s,b)$: machine $\mathcal{M}$ halts on input $s$ with output bit $b$.
  5. (v)
@@
Remark 5.3.2.
As said, certification of a cubic decider inside $\mathsf{RA}$ would not merely refute the decider—it would equip $\mathsf{RA}$ with the means to prove its own consistency. The provability bridge transports $\mathsf{RA}$-proofs through the compiler $f$; the certified decision specification $\mathrm{Dec}_{\mathcal{M}}$ then closes the reflection loop (Figure 2).
Definition 5.3.3 (Internal decision specification).
Let $\operatorname{Sat}_{\Sigma_{1}}(u)$ be the standard $\Sigma_{1}$-satisfaction predicate and $f$ the primitive recursive compiler of Theorem 1.3.1. For a Turing Machine $\mathcal{M}$, define $\mathrm{Dec}_{\mathcal{M}}$ as
$\forall u\;\Big(\operatorname{Sigma1Sent}(u)\rightarrow\big(\operatorname{HaltOut}_{\mathcal{M}}(f(u),1)\leftrightarrow\operatorname{Dioph}_{3}(f(u))\big)\Big).$ (145)
@@
Theorem 5.3.4 (Provable decision on the compiler image yields $\Sigma_{1}$-reflection).
@@
$\mathsf{RA}\vdash\;\forall u\;\Big(\operatorname{Sigma1Sent}(u)\rightarrow\big(\operatorname{Sat}_{\Sigma_{1}}(u)\leftrightarrow\operatorname{Dioph}_{3}(f(u))\big)\Big).$ (146)
@@
By construction. Since $f$ is primitive recursive, its graph is $\mathsf{RA}$-representable (Proposition 1.3.3). Internalizing the compilation step yields a provability bridge: for each $\Sigma_{1}$-sentence code $u$,
@@
$\mathsf{RA}\vdash\;\operatorname{Prov}_{\mathsf{RA}}(u)\rightarrow\operatorname{Prov}_{\mathsf{RA}}\bigl(\ulcorner\operatorname{Dioph}_{3}(f(u))\urcorner\bigr).$ (147)
@@
Combining with $\mathrm{Dec}_{\mathcal{M}}$ and (146) yields $\operatorname{Prov}_{\mathsf{RA}}(u)\rightarrow\operatorname{Sat}_{\Sigma_{1}}(u)$, i.e. $\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$. ∎
Corollary 5.3.5 (Incompleteness Discharge).
@@
Assume $\mathsf{RA}$ is consistent with $\operatorname{Prov}_{\mathsf{RA}}$ satisfying the Hilbert–Bernays–Löb conditions (D1)–(D3) of Theorem 3.5.1. Then no Turing Machine $\mathcal{M}$ satisfies $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$: Theorem 5.3.4 would give $\mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$, hence $\mathsf{RA}\vdash\mathrm{Con}(\mathsf{RA})$—contradicting Gödel’s 2nd [Gödel, 1931; Feferman, 1960; Boolos et al., 2012].
Assuming $\mathrm{Con}(\mathsf{RA})$, no Turing Machine $\mathcal{M}$ satisfies $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$. (If $\mathsf{RA}$ is inconsistent then $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$ holds for all $\mathcal{M}$, so the corollary is stated only under $\mathrm{Con}(\mathsf{RA})$.)